# LATTICES OF FLATS FOR SYMPLECTIC MATROIDS

### v5

@@ 5.8. Lemma:
If the number of atoms in a $C_{n}$ lattice $\mathcal{L}$ is $t$, then the number of atoms of $\left[A,\hat{1}\right]\subseteq\mathcal{L}$ for $A$ an atom of $\mathcal{L}$ is $\leq t-2$.
@@
As $\mathcal{L}$ is atomic, we have for each of its elements a join of atoms. By property (4) of $C_{n}$ lattices the elements that cover $A$ pairwise disjoint (excluding the elements of $A$) and so every element covering $A$ is a choice of some atoms from a set of size $t-2$, namely the set $A(\mathcal{L})\setminus\left\{A,A^{*}\right\}$. There can be at most $t-2$ choices.
To finish our technical preparations we recall the definition of an independent set following Section 4. Using Lemma [5.5], we have for a $C_{n}$ lattice $\mathcal{L}$ of rank $\geq 3$, $A(\mathcal{L})=\mathcal{X}\sqcup\mathcal{X}^{*}$. Therefore, since $\mathcal{L}$ is also atomic it is order isomorphic to a $C_{n}$ lattice on the ground set $A(\mathcal{L})$ sending the atom $A_{i}$ in each element to $i$ and the atom ${A_{i}}^{*}$ to $i^{*}$. In the following theorems, we take independent sets as sets of atoms using the correspondence above. In addition, using lemmas [5.5] and [5.6] we can give a $C_{n}$ lattice structure to the restriction $[B,\hat{1}]$ for every $B\in\mathcal{L}$. As the * function on the restriction is induced by the original function on $\mathcal{L}$, we get the following relation between admissible independent sets:
@@ 5.9. Lemma:
Let $\omega$ be a linear order on the atoms of a $C_{n}$ lattice $\mathcal{L}$ of rank $d\geq 4$ with the first $d-1$ atoms forming a geometric independent set $\mathcal{I}$. Then for every $A_{i}\in A(\mathcal{L})$ there exists a linear order $\omega_{i}$ on the atoms of $\mathcal{L}_{i}=[A_{i},\hat{1}]$ with the atoms that come first in the ordering being those that cover some $A_{j}$ where $j<i$ and the first $d-2$ atoms form a geometric independent set of $\mathcal{L}_{i}$. Moreover, if $\omega$ has that no atom is the star of the previous atom we can construct $\omega_{i}$ also having this property. We denote such an order $\omega$ as an admissible ordering.
@@
    if $i\leq d-1$, then using Theorem [4.9], the first $d-2$ atoms of $\omega_{i}$ will be $A_{1}\vee A_{i},...,A_{d-1}\vee A_{i}$ and so a geometric independent set of $\mathcal{L}_{i}$ and $\omega_{i}$ the required order.
@@
    If $i\geq d$, let $\mathcal{I}^{\prime}$ be an extension of $A_{i}$ to a geometric independent set of size $d-1$ contained in $\mathcal{I}\cup\left\{A_{i}\right\}$. We remark that if ${A_{i}}^{*}\in\mathcal{I}$ then ($(\mathcal{I}\setminus\left\{{A_{i}}^{*}\right\})\cup\left\{A_{i}\right\}$ is a geometric independent set according to the lemma [4.7]. Otherwise, we have one of $A_{i},{A_{i}}^{*}$ not in $\vee\mathcal{I}$ and by Theorem [4.9] there exist some $j\leq d-2$ such that $A_{i},{A_{i}}^{*}\notin\vee(\mathcal{I}\setminus A_{j})$. Therefore, ($(\mathcal{I}\setminus\left\{A_{j}\right\})\cup\left\{A_{i}\right\}$ is a geometric independent set. Let $\omega^{\prime}$ be the order on $A(\mathcal{L})$ obtained from $\omega$ by moving $A_{i},A_{j}$ between $A_{d-1}$ and $A_{d}$. We then have $\omega^{\prime}_{i}$, the induced order by $\omega^{\prime}$ on $A(\mathcal{L}_{i})$ being the required order.
As no non-maximal geometric independent set contains an atom and its star, we can extend $\mathcal{I}$ to a linear order $\omega$ with $A_{i}\neq{A^{*}_{i+1}}$ for every $i\leq\left|A\left(\mathcal{L}\right)\right|$. Notice that we cannot always extend the $d-2$ admissible independent set constructed above in $\mathcal{L}_{i}$ in the way specified. However, since we can always rearrange the atoms of the independent set (it always contains more than two elements), a problem occurs only if the atoms of $\mathcal{L}_{d+1}$ covering $A_{d+1}$ are exactly our independent set + an extra atom and its star. In this case, we have $\left\{A_{d+1},A_{j}\right\}$ (as above) an independent set and we can change $\mathcal{I}^{\prime}$ so that it contains $\left\{A_{i},A_{j}\right\}$. As there is only one star of $A_{d}\vee A_{d+1}$ in $\mathcal{L}_{d+1}$, this solves our problem.
@@ 5.10. Corollary:
If $\left\{a_{1},...,a_{k}\right\}$ is a geometric independent set of a graded, bounded, and atomic lattice $\mathcal{S}$, then $\left\{a_{1}\vee a_{i},...,a_{k}\vee a_{i}\right\}\setminus\left\{a_{i}\right\}$ is a geometric independent set of $\left[a_{i},\hat{1}\right]$.
@@ 5.11. Theorem:
First notice that the conditions for a recursive atom ordering are met if the rank of $\mathcal{L}$ is $2$, since then all atoms are covered by $\hat{1}$. Another separated case is $rank(\mathcal{L})=3$, in this case taking any linear order $\omega$ on $A(\mathcal{L})$ with $A_{i}\neq{A_{i+1}}^{*}$ for every $i\leq\left|A\left(\mathcal{L}\right)\right|$ is a recursive atom ordering. The first condition is obvious, as all the elements of rank $2$ are covered by $\hat{1}$. The second condition is true, as every two consecutive atoms are covered by a rank $2$ element and the only two elements not covered by a rank $2$ element are stars of each other for which their join is $\hat{1}$.
@@
We now prove that any linear order on the atoms of a $C_{n}$ lattice $\mathcal{L}$ of rank $d$ with its first $d-1$ atoms forming a geometric independent set and no atom being the star of the previous atom is a recursive atom ordering. We assume that the statement is true for every $C_{n}$ lattice $\mathcal{L}^{\prime}$ with $\leq t-2$ atoms and prove it for a $C_{n}$ lattice $\mathcal{L}$ with $t$ atoms.
Let $A_{1},...,A_{t}$ be such an atom ordering. For the first condition, we just apply lemma [5.8] and lemma [5.7] to ensure that $\left[A_{i},\hat{1}\right]\subseteq\mathcal{L}$ has $\leq t-2$ atoms. Therefore, if $rank(\mathcal{L})\geq 5$ we get a recursive atom ordering using the induction hypothesis, and otherwise we get a recursive atom ordering by our separate cases.
The proof of the second condition is the same as in the separate case $rank(\mathcal{L})=3$ as by Lemma [5.8] no atom is the star of the previous atom in linear order $\omega_{i}$.
@@
In this section we prove a characterization of $C_{n}$ lattices by recursive atom ordering inspired by the work in [8], our goal is to give a minimal list of orders that are always and never recursive atom orderings equivalent to being a $C_{n}$ lattice. We start with a definition of strongly admissible sets characterizing non maximal geometric independent sets in $C_{n}$ lattices.
@@ 6.1. Definition:
A set $A$ of atoms will be called strongly admissible if it is obtained from a size $\left|A\right|-1$ strongly admissible set $B$ with $a^{*}\nless\vee B$ for $a\in A\setminus B$.
We observe that every strongly admissible set is indeed admissible and that for any admissible element $X$ of a bounded, atomic, and graded lattice $\mathcal{L}$ we have a spanning strongly admissible independent set. Next, we present a characterization of $C_{n}$ lattices similar to the characterization of geometric lattices given in Theorem [3.9].
@@ 6.2. Theorem:
A bounded, atomic, and graded lattice $\mathcal{L}$ is order isomorphic to a $C_{n}$ lattice iff there is a partition of the atoms of $\mathcal{L}$ to pairs ${1},{1}^{*},{2},{2}^{*},...,{k},{k}^{*}$ such that the non-maximal geometric independent sets are exactly the strongly admissible non maximal independent sets.
@@
If $\mathcal{L}$ is a $C_{n}$ lattice, then equality is a consequence of Lemma [4.7]. In the other direction, every atomic lattice $\left(\mathcal{L},\leq\right)$ is order isomorphic to a set of subsets, ordered by inclusion by defining $x=\left\{a\in A(\mathcal{L})\mid a\leq x\right\}$ for every $x\in\mathcal{L}$. We prove the four conditions in Definition [2.1]:
@@
    As $\mathcal{L}$ is bounded, we just have to denote the minimal set $\phi$ and the maximal set $J$. Notice that since $\mathcal{L}$ is atomic, we must have $J=\left[k\right]\sqcup\left[k\right]^{*}$.
@@
    Let $A\in\mathcal{L}$ be an inadmissible set, $\left\{i,i^{*}\right\}\subset A$. As $\left\{i,i^{*}\right\}$ is an independent set we can extend it to a size $rank(A)$ independent set $I$ that spans $A$ (this extension only exists for sizes $\leq 2$ sets of atoms). As $I$ is inadmissible and geometric, we must have $rank(I)=rank(A)=rank(\mathcal{L})$ and so $A=J$.
@@
    This property is always true for an atomic and bounded lattice of subsets. $A\wedge B\subseteq A\cap B$ and if there is an atom $i\in A\cap B$ not in the meet then $A\wedge B\subsetneq(A\wedge B)\vee\left\{i\right\}\subset A,B$ which is a contradiction to the definition of $A\wedge B$, therefore $A\wedge B=A\cap B$.
@@
    Let $B,C\in\mathcal{L}$ be two elements that cover some $A\in\mathcal{L}$, since $B\cap C\in\mathcal{L}$ we must have $B\cap C=A$. Furthermore, if $A$ is not covered by $J$, then for $i\in A$ we have $I\vee i^{*}$ not geometric for every spanning geometric set of $A$ containing $i$. Therefore, there is no element covering $A$ and containing $i$. It remains to prove that if $i\in J\setminus(A\sqcup A^{*})$ then there exists $B\in\mathcal{L}$ covering $A$ with $i\in B$. Take a strongly admissible independent set $I$ that spans $A$, making $I\cup\left\{i\right\}$ also a strongly admissible independent set. Therefore, $I\cup\left\{i\right\}$ is geometric and we have $rank(I\cup\left\{i\right\})=rank\left(A\right)+1$ and the element of $\mathcal{L}$ spanned by $I\cup\left\{i\right\}$ is the desired $B$.
@@ 6.3. Corollary:
A bounded, atomic, and graded lattice $\mathcal{L}$ is order isomorphic to a $C_{n}$ lattice iff there is a partition of the atoms of $\mathcal{L}$ into pairs ${1},{1}^{*},{2},{2}^{*},...,{k},{k}^{*}$ such that the non-maximal geometric independent sets are admissible and strongly admissible independent sets are geometric.
@@ 6.4. Theorem:
A bounded, atomic and graded lattice $\mathcal{L}$ of rank $d\lneq k=\left|A\left(\mathcal{L}\right)\right|$ is order isomorphic to a $C_{n}$ lattice iff there is a partition of the atoms of $\mathcal{L}$ to pairs ${1},{1}^{*},{2},{2}^{*},...,{2k},{2k}^{*}$ such that the following conditions hold:
  1. (1)
    Every strongly admissible independent set of size $\leq d-1$ (in any order) can be extended to a recursive ordering.
  2. (2)
    A linear order $a_{1},...,a_{2k}$ starting with a geometric independent set of size $d-1$ is a recursive atom ordering iff $a_{i}\neq{a_{i+1}}^{*}$ for $i\in[2k-1]$.
@@
If $\mathcal{L}$ is a $C_{n}$ lattice, every strongly admissible independent set of size $\leq d-1$ is geometric, and we have proven in Theorem [5.10] that such recursive atom orderings exist.
For the other direction, we use corollary [6.3]. Throughout the proof, we assume that $d\geq 4$ as the cases $d=2,3$ are obvious. Every such lattice of rank $2$ is a $C_{n}$ lattice, and for $d=3$ we only need to check that admissible pairs are of rank $2$ and inadmissible pairs are of rank $3$.
We first observe that every independent set $I$ for which every linear order of its elements can be extended to a recursive atom ordering is geometric. Therefore, every strongly admissible independent set of size $\leq d-1$ is geometric.
We proceed to prove that no element of the form $i\vee i^{*}$ is of rank $2$. Assume there is such an $i$ and notice that if $rank(i\vee i^{*})\lneq d$ we can always extend $\left\{i,i^{*}\right\}$ to a size $d-1$ geometric independent set $I$. Therefore, we can construct two recursive atom ordering of the form $i^{*},j,i,\omega$ and $i,j,i^{*},\omega$, with $j\in I$ and $\omega$ being a linear ordering of $A(\mathcal{L})\setminus\left\{i,i^{*},j\right\}$. We will now prove that if $rank(i\vee i^{*})=2$ then $j,i,i^{*},\omega$ is also a recursive atom ordering. Using the two previously constructed orderings, it remains to show that $\left[i,\hat{1}\right]$ has a recursive atom ordering starting with $i\vee j$. If $d\geq 5$ we can take a recursive atom ordering on $\mathcal{L}$ starting with $j,i,t,i^{*}$ for some $t\in I$ to obtain a recursive atom ordering of $\left[i,\hat{1}\right]$ starting with $i\vee j$. Otherwise we need to consider the case $d=4$, it is enough to find some atom $t$ with $\left\{i,j,t\right\}$ a strongly admissible independent set. This set exists unless $i\vee j=\vee T$ with $T$ a transversal. Using a symmetry argument, we must also have $i^{*}\vee j=\vee T$, which results in $A(\mathcal{L})=\left\{i,i^{*},j,j^{*}\right\}$ contradicting $rank(\mathcal{L})\lneq\left|A\left(\mathcal{L}\right)\right|$.
We now know that no element of the form $i\vee i^{*}$ is of rank $2$ and using the recursive atom ordering $i,j,i^{*},\omega$ we get $j\leq i\vee i^{*}$. As this is true for every $j\in I$, we must have $rank(i\vee i^{*})\geq d-1$. Let $j\nleq i\vee i^{*}$, we have $\left\{i,j\right\}$ a strongly geometric independent set and so of rank $2$. Therefore, we can extend $\left\{i,j\right\}$ to a geometric independent set of rank $d-1$, either containing $j$ or not, in both cases we can construct a recursive atom ordering containing the sequence $i,j,i^{*}$. Consequently $rank(i\vee i^{*})=d$ and there does not exist a non-maximal inadmissible independent set.

### v6

@@ 5.8. Lemma:
If the number of atoms in a $C_{n}$ lattice $\mathcal{L}$ is $t$, then the number of atoms of $\left[A,J\right]\subseteq\mathcal{L}$ for $A$ an atom of $\mathcal{L}$ is $\leq t-2$.
@@
As $\mathcal{L}$ is atomic, we have that each of its elements is a join of some of its atoms. By property (4) of $C_{n}$ lattices the elements that cover $A$ pairwise disjoint (excluding the elements of $A$) and so every element covering $A$ is a choice of some atoms from a set of size t2t-2, namely the set $A(\mathcal{L})\setminus\left\{A,A^{*}\right\}$. There can be at most t2t-2 choices.
To finish our technical preparations, we observe that we can always work with $C_{n}$ lattices in which every atom is a singleton. This is done for geometric lattices by deleting all the elements in an atom except one, the process is discussed in [1]. For $C_{n}$ lattices $\mathcal{L}$ of rank $\geq 3$, we have seen that $A(\mathcal{L})=\sqcup_{i=1}^{k}(A_{i}\sqcup A_{i}^{*})$. Therefore, we can use the same deletion process, deleting all pairs $a,a^{*}$ in every pair of atoms $A_{i},A_{i}^{*}$ except one.
In the following Theorems, we take independent sets as sets of atoms using the process above. In addition, using Lemmas 5.6 and 5.7 we can give a $C_{n}$ lattice structure to the restriction $[A,J]$ for every $A\in A(\mathcal{L})$. As the * function on the restriction is induced by the original function on $\mathcal{L}$, we get the following relation between admissible independent sets:
@@ 5.9. Lemma:
Let $\omega$ be a linear order on the atoms of a $C_{n}$ lattice $\mathcal{L}$ of rank $d\geq 4$ with the first d1d-1 atoms forming a geometric independent set $\mathcal{I}$. Then for every $A_{i}\in A(\mathcal{L})$ there exists a linear order $\omega_{i}$ on the atoms of $\mathcal{L}_{i}=[A_{i},\hat{1}$] with the atoms that come first in the ordering being those that cover some $A_{j}$ where $j<i$ and the first d2d-2 atoms form a geometric independent set of $\mathcal{L}_{i}$.
@@
    if $i\leq d-1$, then using Theorem 4.9, the first d2d-2 atoms of $\omega_{i}$ will be $A_{1}\vee A_{i},...,A_{d-1}\vee A_{i}$. This is a geometric independent set of $\mathcal{L}_{i}$, and therefore $\omega_{i}$ is the required order.
@@
    If $i\geq d$, we again have two options. If $A_{i}^{*}\notin\vee\mathcal{I}$ then, using Theorem 4.11 there exist $\mathcal{I}^{\prime}\subsetneq\mathcal{I}$ such that $\mathcal{I}^{\prime}\cup\left\{A_{i}\right\}$ is a geometric independent set of size d1d-1. Otherwise, $A_{i}^{*}\in\vee\mathcal{I}$ and by Theorem 4.11 there exist $\mathcal{I}^{\prime}\subsetneq\mathcal{I}$ such that $\mathcal{I}^{\prime}\cup\left\{A_{i}^{*}\right\}$ is a geometric independent set of size d1d-1. applying Lemma 4.10 we have $\mathcal{I}^{\prime}\cup\left\{A_{i}\right\}$ also a geometric independent set of size d1d-1.
    In both cases we have any order $\omega_{i}^{\prime}$ starting with $\left\{A_{i}\vee A_{j}\mid A_{j}\in\mathcal{I}^{\prime}\right\}$ the required order.
@@ 5.10. Corollary:
If $\left\{a_{1},...,a_{k}\right\}$ is a geometric independent set of a finite, graded, and atomic lattice $\mathcal{S}$, then $\left\{a_{1}\vee a_{i},...,a_{k}\vee a_{i}\right\}\setminus\left\{a_{i}\right\}$ is a geometric independent set of $\left[a_{i},\hat{1}\right$].
@@ 5.11. Theorem:
First, notice that the conditions for a recursive atom ordering are met if the rank of $\mathcal{L}$ is $2$, since then all atoms are covered by $J$. Another separated case is $rank(\mathcal{L})=3$, in this case taking any linear order $\omega$ on $A(\mathcal{L})$ with $A_{1}\neq{A_{2}}^{*}$ is a recursive atom ordering. The first condition is obvious, as all the elements of rank $2$ are covered by $J$. To see the second condition holds, observe that every pair of atoms which are not stars of each other is covered by a rank $2$. If a pair of atoms $A_{i},A_{j}$ with $i\lneq j$ are stars of each other, then $\left\{i,j\right\}\neq\left\{1,2\right\}$. Therefore, $A_{1}\vee A_{j}$ covers $A_{1},A,j$ or $A_{2}\vee A_{j}$ covers $A_{2},A,j$.
@@
We now prove that any linear order on the atoms of a $C_{n}$ lattice $\mathcal{L}$ of rank $d$ with its first d1d-1 atoms forming a geometric independent set is a recursive atom ordering. We assume that the statement is true for every $C_{n}$ lattice $\mathcal{L}^{\prime}$ with $\leq t-2$ atoms and prove it for a $C_{n}$ lattice $\mathcal{L}$ with $t$ atoms.
Let $A_{1},...,A_{t}$ be such an atom ordering. For the first condition, we just apply Lemma 5.8 and Lemma 5.9 to ensure that $\left[A_{i},J\right]\subseteq\mathcal{L}$ has $\leq t-2$ atoms. Therefore, we get a recursive atom ordering using the induction hypothesis.
The proof of the second condition is the same as in the separate case $rank(\mathcal{L})=3$ as linear order $\omega_{i}$ starts with a geometric independent set which cannot contain an atom and its star..
@@
In this section we prove a characterization of $C_{n}$ lattices by recursive atom ordering inspired by the work in [8], our goal is to give a minimal list of orders that are always recursive atom orderings equivalent to being a $C_{n}$ lattice. We start with a definition of strongly admissible sets characterizing non maximal geometric independent sets in $C_{n}$ lattices.
@@ 6.1. Definition:
A set $A$ of atoms is said to be strongly admissible if either $A=\emptyset$, or there exists a strongly admissible subset $B\subsetneq A$ with $\left|B\right|=\left|A\right|-1$ such that for the unique element $a\in A\setminus B$, we have $a^{*}\nless\vee B$.
We observe that every strongly admissible set is indeed admissible and that for any admissible element $X$ of a finite, atomic, and graded lattice $\mathcal{L}$ we have a spanning strongly admissible independent set. Next, we present a characterization of $C_{n}$ lattices similar to the characterization of geometric lattices given in Theorem 3.9.
6.2. Lemma:
Let $\mathcal{L}$ be a finite, atomic, and graded lattice. For every $a,b\in A(\mathcal{L})$ There exists a geometric independent set $a,b\in A$ such that $rank(a\vee b)=|I|$ .
Proof:
Using Lemma 3.5 we will construct a spanning geometric independent set for $a\vee b$, $B=\left\{a=a_{1},a_{2},...,a_{rank(a\vee b)}\right\}$ such that $n=rank(\vee_{i=1}^{n}a_{i})$ for all $1\leq n\leq rank(a\vee b)$ and $\vee B=a\vee b$. Let $1\leq n\leq rank(a\vee b)$ be minimal so that $b\in\vee_{i=1}^{n}a_{i}$ and define $I=(B\setminus\left\{a_{n}\right\})\cup\left\{b\right\}$. We again have by Lemma 3.5 that $I$ is independent and so our desired geometric independent set.
@@ 6.3. Theorem:
A finite, atomic, and graded lattice $\mathcal{L}$ is order isomorphic to a $C_{n}$ lattice if and only if there is a partition of the atoms of $\mathcal{L}$ to pairs ${1},{1}^{*},{2},{2}^{*},...,{k},{k}^{*}$ such that the non-maximal geometric independent sets are exactly the strongly admissible non maximal sets.
@@
If $\mathcal{L}$ is a $C_{n}$ lattice, then equality is a consequence of Lemma 4.8. Conversely, every atomic lattice $\left(\mathcal{L},\leq\right)$ is order isomorphic to a set of subsets, ordered by inclusion by defining $x=\left\{a\in A(\mathcal{L})\mid a\leq x\right\}$ for every $x\in\mathcal{L}$. We prove the four conditions in Definition 2.3:
@@
    As $\mathcal{L}$ is a finite lattice, we just have to denote the minimal set $\phi$ and the maximal set $J$. Notice that since $\mathcal{L}$ is atomic, we must have $J=\left[k\right]\sqcup\left[k\right]^{*}$.
@@
    Let $A\in\mathcal{L}$ be an inadmissible set, $\left\{i,i^{*}\right\}\subset A$. Using Lemma 6.2 we can extend $\left\{i,i^{*}\right\}$ to a size $rank(A)$ independent set $I$ that spans $A$. As $I$ is inadmissible and geometric, we must have $rank(I)=rank(A)=rank(\mathcal{L})$ and so $A=J$.
@@
    This property is always true for a finite atomic lattice of subsets. $A\wedge B\subseteq A\cap B$ and if there is an atom $i\in A\cap B$ not in the meet then $A\wedge B\subsetneq(A\wedge B)\vee\left\{i\right\}\subset A,B$ which is a contradiction to the definition of $A\wedge B$, therefore $A\wedge B=A\cap B$.
@@
    Let $B,C\in\mathcal{L}$ be two elements that cover some $A\in\mathcal{L}$, since $B\cap C\in\mathcal{L}$ we must have $B\cap C=A$. Furthermore, if $A$ is not covered by $J$, then for $i\in A$ we have $I\vee i^{*}$ not geometric for every spanning geometric set $I$ of $A$ . Therefore, there is no element covering $A$ and containing $i^{*}$. It remains to prove that if $j\in J\setminus(A\sqcup A^{*})$ then there exists $B\in\mathcal{L}$ covering $A$ with $j\in B$. Take a strongly admissible independent set $I$ that spans $A$, making $I\cup\left\{i\right\}$ also a strongly admissible independent set. Therefore, $I\cup\left\{i\right\}$ is geometric and we have $rank(I\cup\left\{i\right\})=rank\left(A\right)+1$. The element of $\mathcal{L}$ spanned by $I\cup\left\{i\right\}$ is the desired $B$.
@@ 6.4. Corollary:
A finite, atomic, and graded lattice $\mathcal{L}$ of rank $d$ is order isomorphic to a $C_{n}$ lattice if and only if there is a partition of the atoms of $\mathcal{L}$ to pairs ${1},{1}^{*},{2},{2}^{*},...,{k},{k}^{*}$ such that the size d1d-1 geometric independent sets are exactly the size d1d-1 strongly admissible sets.
6.5. Lemma:
Let $\mathcal{L}$ be a finite, atomic, and graded lattice, and let $I$ be a set of atoms of $\mathcal{L}$. If every ordering of the atoms of $\mathcal{L}$ starting with $I$ yields a recursive atom ordering, then $I$ is a geometric independent set.
Proof:
We prove this by induction on the size of II. The base case $|I|=1$ is trivially geometric, since the rank of a single atom is 1.
Assume that every proper subset of $I=\{a_{1},a_{2},\dots,a_{|I|}\}$ is geometric. We aim to show that $I$ itself is geometric. Let $\omega_{1}$ be a recursive atom ordering of $\mathcal{L}$ that begins with the atoms of $I$, i.e.,
$\omega_{1}=a_{1}\lneq a_{2}\lneq\cdots\lneq a_{|I|}\lneq\cdots$
Now consider the induced recursive atom ordering $\omega_{2}$ on the interval $a_{|I|},\hat{1}$]. This ordering begins with the set:
$I^{\prime}=\{a_{1}\vee a_{|I|},a_{2}\vee a_{|I|},\dots,a_{|I|-1}\vee a_{|I|}\}$
According to the induction hypothesis, each pair $\{a_{i},a_{|I|}\}$ is geometric, so each join $a_{i}\vee a_{|I|}$ has rank 2. Moreover, these joins are distinct, since the join of three atoms is of rank $3$. We repeat this process iteratively: for each $i=3,\dots,|I|-1$, define $\omega_{i}$ as the recursive atom ordering on $b_{|I|+1-i},\hat{1}$], where $b_{|I|-i}$ is the (|I|i)(|I|-i)th atom in $\omega_{i-1}$. The order $\omega_{|I|-1}$ begins with two elements of the form:
$\left\{(\vee_{\underset{i\neq j,k}{1\leq i\leq\left|I\right|}}a_{i})\vee a_{j},(\vee_{\underset{i\neq j,l}{1\leq i\leq\left|I\right|}}a_{i})\vee a_{j}\right\}$
for distinct $j,k,l\in I$.
Since $\omega_{|I|-1}$ is a recursive atom ordering, the join of these two elements must have rank $2$ in the interval $\vee_{\underset{i\neq l,k}{1\leq i\leq\left|I\right|}}a_{i},\hat{1}$]. Thus, the full join satisfies:
$rank(\vee_{1\leq i\leq\left|I\right|}a_{i})=rank(((\vee_{\underset{i\neq j,k}{1\leq i\leq\left|I\right|}}a_{i})\vee a_{j})\vee((\vee_{\underset{i\neq j,l}{1\leq i\leq\left|I\right|}}a_{i})\vee a_{j}))$
$rank((\vee_{\underset{i\neq j,k}{1\leq i\leq\left|I\right|}}a_{i})\vee a_{j})+1=|I|$
Hence, I is a geometric independent set.
@@ 6.6. Theorem:
Let $\mathcal{L}$ be a finite, atomic, and graded lattice of rank $3\leq d\lneq 2k=\left|A\left(\mathcal{L}\right)\right|$. then $\mathcal{L}$ is order isomorphic to a $C_{n}$ lattice if and only if the following condition holds:
There exists a partition of the atom set $A(\mathcal{L})$ into $k$ disjoint pairs:
$\left\{{1},{1}^{*}\right\},\left\{{2},{2}^{*}\right\},...,\left\{{k},{k}^{*}\right\}$
such that every linear ordering
$a_{1}\lneq...\lneq a_{2k}$
of the atoms beginning with a subset of size d1d-1 that is either strongly admissible or geometric independent, constitutes a recursive atom ordering if and only if the condition
$a_{i}\neq{a_{i+1}}^{*}\textbf{ for all }i\in[d-2]$
is satisfied.
@@
If $\mathcal{L}$ is a $C_{n}$ lattice, every strongly admissible independent set of size $\leq d-1$ is geometric, and we have proven in Theorem 5.11 that such a recursive atom ordering exists.
To prove the converse, we invoke Corollary 6.4. Throughout the proof, we assume $d\geq 4$, since the case $d=3$ is straightforward. When $d=3$, condition (1) of Definition 5.3 holds trivially, and condition (2) immediately implies that admissible pairs of atoms have rank $2$, while inadmissible pairs have rank $3$.
Lemma 6.5 states that if every ordering of the atoms of $\mathcal{L}$ starting with a set $I$ yields a recursive atom ordering, then I is a geometric independent set. Therefore, every strongly admissible set is a geometric independent set.
We now aim to prove that $rank(i\vee i^{*})=d$ for every $i\in[k]$. This result implies that any geometric independent set of size less than $d$ must be strongly admissible. Suppose, for contradiction, that this is not the case. Then, by Lemma 6.2, there exists a geometric independent set $A\subseteq A(\mathcal{L})$ of size d1d-1 that includes both $i$ and $i^{*}$. Under this assumption, we obtain a recursive atom ordering $\omega$ of the form:
$i\prec j\prec i^{*}\prec\cdots$
for some $j\in A$. It is straightforward to verify that the modified ordering $\omega_{i,i^{*}}$, obtained by swapping the positions of $i$ and $i^{*}$, also satisfies the conditions of a recursive atom ordering.
We now verify that condition (1) of Definition 5.3 holds for the ordering $\omega_{i,j}$, which is obtained from a recursive atom ordering $\omega$ by swapping the positions of $i$ and $j$. Since $\omega$ is a recursive atom ordering, it suffices to show that the interval $i,\hat{1}$] admits a recursive atom ordering that begins with $i\vee j$.
  • Case 1: $d\geq 5$
    In this case, we can construct a recursive atom ordering of $\mathcal{L}$ that begins with the sequence $j\prec i\prec t\prec i^{*}$ for some atom $t\in I$. This ordering induces a recursive atom ordering on $i,\hat{1}$] that starts with $i\vee j$, as required.
  • Case 2: $d=4$
    Here, it suffices to find an atom $t$ such that $\{i,j,t\}$ forms a strongly admissible independent set. Such a set exists unless $i\vee j=\bigvee T$ for some transversal $T$. However, by symmetry, this would also imply $i^{*}\vee j=\bigvee T$, leading to the conclusion that the atom set is $A(\mathcal{L})=\{i,i^{*},j,j^{*}\}$ which contradicts the assumption that $rank(\mathcal{L})<|A(\mathcal{L})|$.
Since condition (1) of Definition 5.3 is satisfied for the ordering $\omega_{i,j}$, yet $\omega_{i,j}$ fails to be a recursive atom ordering, it follows that condition (2) must be violated. Consequently, we deduce that:
[…]
This contradicts condition (1) of Definition 5.3 as applied to the ordering $\omega$.